OpenStudy (anonymous):

Jill wants to solve the following system using the elimination method: y = x + 8 y = 3x − 11 What number should the equation y = x + 8 be multiplied by to eliminate x? −1 −2 −3 3

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

I am thinking -3

OpenStudy (anonymous):

am I correct

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

yeah because -3x + 3x = 0

OpenStudy (anonymous):

It depends on what elimination method

OpenStudy (anonymous):

thanks I gave a medal can you medal back??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

If addition, -3, if substitution 3

OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system given below? x − 4y = 7 5x + 9y = 6 (3, −1) (3, 1) (1, −3) (−1, −3)

OpenStudy (anonymous):

@AliSurfCat What was the elimination method for the first question?

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

x = 7 + 4y so insert 7 + 4y for x in the second equation 5x + 9y = 6

OpenStudy (anonymous):

It must have been included in the question. They wouldn't have said "using the elimination method:" without including a specific elimination method

OpenStudy (anonymous):

Katross what is the answer?

OpenStudy (anonymous):

You should try to figure it out yourself. Don't just ask for the answer.

OpenStudy (anonymous):

I have figured them all out without no help but this one I dont know

OpenStudy (anonymous):

5(7+4y) +9y = 6 solve from here

OpenStudy (anonymous):

it is (3,-1)

OpenStudy (anonymous):

yep